### v2

# Mapping the Unseen: Unified Promptable Panoptic Mapping with Dynamic Labeling using Foundation Models

Abstract
In robotics and computer vision, semantic mapping remains an important challenge due to the increasing demand for machines that can comprehend and interact with complex environments. However, conventional panoptic mapping approaches are constrained by preset semantic classes, which makes them inefficient when dealing with novel or unforeseen objects. In response to this constraint, we propose the Unified Promptable Panoptic Mapping (UPPM) approach. UPPM leverages the latest advancements in foundation models to generate labels on demand, without additional training or fine-tuning. By integrating a dynamic labeling strategy into conventional panoptic mapping methods, UPPM significantly improves adaptation and versatility while maintaining desirable map reconstruction performance. We demonstrate our approach on real-world and simulated datasets. Results show that UPPM can accurately reconstruct scenes and segment objects while generating rich semantic labels through natural language interactions. A series of ablation studies validated the advantages and the limitations, of foundation-model-based labeling over fixed label sets.
keywords:
Semantic Scene Understanding , Deep Learning for Visual Perception , Mapping
journal: Pattern Recognition Letters
[…]
Center for AI Technology, Skolkovo Institute of Science and Technology, addressline=, city=Skolkovo, postcode=121205, state=Moscow, country=Russia
1 Introduction
Panoptic mapping is a key element in enabling robots to perceive and reconstruct their surroundings with better comprehension. The representation of rich semantic information forms the building block of intelligent machine perception [1]. However, existing methods [2, 3, 4] have issues due to their reliance on rigid predefined class sets, which limits their flexibility to unforeseen objects or dynamic environments. To be applicable in various real-world scenarios, these systems require either a large amount of labeled data [5, 6] or controlled environmental conditions. Integrating open-set techniques [7, 8] is a potential way to improve robustness and generalizability.
In this paper, we present a prompt-based panoptic mapping pipeline that uses on-demand label generation based on natural language inputs. As a result of this approach, robots can dynamically acquire and apply object labels, allowing deeper semantic understanding and enhanced flexibility. We define Dynamic Labeling as the process of assigning semantically unified categories (see fig. 3) to detected objects in previously unseen environments, all while preserving the rich labels generated through open-vocabulary methods. Thorough evaluations and ablation studies on real-world and simulated datasets show that UPPM scene reconstruction is accurate and the dynamic label generation works well. This study opens the door to more natural human-robot communication and flexible machine perception in dynamic contexts. (Webpage)
2 Related Work
Semantic mapping and visual SLAM have been active research areas, with methods proposed for dense semantic mapping [9, 10], object-centric mapping [4, 3], as well as keypoint-based object-level SLAM [11]. These methods have contributed significantly to understanding complex real-world scenarios by focusing on object-level semantic mapping. SemanticFusion [9] and DA-RNN [10] are both dense semantic mapping approaches that assign semantic labels to map elements such as voxels and surfels. While such methods allow for a comprehensive scene understanding, they struggle with recognizing individual objects within the scene. In contrast, object-centric methods such as SLAM++ [4] and Fusion++ [3] have shown a strong focus on reconstructing particular objects, but they often lack the ability to incorporate the semantics and geometry of background regions, limiting their capacity for holistic understanding of the scene.
Furthermore, the recent advances in panoptic mapping have tried to overcome the limitations of prior approaches. Panoptic Multi-TSDFs [12] and Panoptic Fusion [2] have made significant contributions to the field by enabling flexible representations for online multi-resolution volumetric mapping, with a focus on long-term dynamic scene consistency and semantic understanding at different levels of granularity. These panoptic techniques can capture both semantic and geometric information, resulting in a more complete understanding of scenes. In spite of this, these methods have a limited ability to adapt to novel objects and environments because they rely on fixed sets of semantic classes. Our framework aims to overcome these restrictions by utilizing foundation models for dynamic label generation.
The proposed UPPM is inspired by the recent breakthroughs of generative AI and foundation models [13, 14, 15, 16] in a wide range of tasks, such as image captioning [17, 18], object detection [19, 20], image segmentation [13], event understanding [21], etc.
UPPM provides a more flexible and versatile labeling technique by allowing on-demand, dynamic label generation for mapped objects using natural language prompts. Compared to point-level approaches [22, 23], UPPM, based on [12], is a more effective method for assigning semantics at the object level. This makes our dynamically-labeled maps more efficient allowing UPPM to be used in a wide range of downstream tasks such as localization, as shown in fig. 1, and navigation, facilitating interactive robotic perception and human-robot communication.
Figure 1: Our UPPM enables interactive scene exploration and object retrieval using natural language prompts, employing query postprocessing and semantic textual similarity for enhanced accuracy. By displaying an environment with a table and dynamic labels attached to it, the application demonstrates the UPPM reconstructed map. Dynamic labels enable the system to respond to four distinct user prompts by pointing to the same object.
3 Methodology
UPPM addresses the visual semantic mapping problem by utilizing foundation models for dynamic label generation. Our approach uses posed RGBD data to generate panoptic segmentation, which serves as the foundation for reconstructing the 3D panoptic volumetric map. This section describes our method, with the mapping pipeline shown in fig. 2.
Figure 2: System Overview of Unified Promptable Panoptic Mapping (UPPM). The UPPM pipeline requires posed RGBD images. VLFE++ generates an open-vocabulary set of classes based on RGB inputs. For each class in the open-vocabulary set, STS predicts its size and parent class. This data is then fed into the Open-set Object Detection & Segmentation modules, where detection is refined by NMS to ensure enhanced accuracy. Combining outputs from previous steps with RGBD inputs, we adapt the panoptic mapping approach to accept dynamic labels (i.e. unified semantics), ensuring compatibility without overstating deviations. The final output is a dynamically labeled multi-resolution multi-TSDF map.
3.1 VLFE++
The input of this stage is the RGB images from the dataset used. As shown in fig. 2, the VLFE++ is composed of interconnected components which generate open-set classes from an RGB image, as described below:
  1. 1.
    Visual Linguistic Features Extraction (VLFE): Using a pre-trained Visual Language Foundation Model, Tag2Text [18], the initial descriptions of the environment will be generated. Tag2Text encodes input images to extract both visual and semantic information, producing caption and identified tags that will be used in the following steps.
  2. 2.
    Part-Of-Speech (POS) Tagging: Using an average perceptron network [24], we detect probable objects and important attributes in the scene by grammatically labeling the generated textual description. Our primary focus here is on extracting nouns and noun phrases, as well as any associated adjectives, in order to provide rich contextual information about the detected objects in the image.
  3. 3.
    Lemmatization: After part-of-speech tagging, lemmatization [25] occurs, which converts inflected or derived word forms into their base or dictionary form (lemma). For instance, ”apples” becoming ”apple” and ”chairs” transforming to ”chair”. This produces normalized representations of words, which then simplify the extracted keywords while keeping their fundamental meanings.
At the end of this stage, the Open-vocabulary set of classes for the given RGB Images will be ready to be used as input for the next component in the pipeline, namely STS.
3.2 Semantic-Textual Similarity (STS)
The input of this stage is the Open-vocabulary set of classes that has been generated from the previous component, VLFE++.
Semantic-Textual Similarity (STS) quantifies the similarity between two pieces of text by considering their meaning. In the context of UPPM, we apply STS, as shown in fig. 2, to predict the size and parent class (corresponding COCO-Stuff class [26]) for each class in the open-vocabulary set. In other words, We define $C$ as the set of COCO-Stuff [26] classes, where each class $(c_{i}\in C)$ has a corresponding size attribute $(s_{i})$, which takes one of three values - small ($s=1$), medium ($s=2$), or large ($s=3$). Our goal is to determine the most appropriate COCO-Stuff class $\hat{c}$ for any given input class $c$:
$\hat{c}=\operatorname*{arg\,max}_{c’\in C}sim(E(c),E(c’))$ ’ ) ) (1)
where $sim(\cdot)$ denotes a similarity function between embeddings and $E(\cdot)$ represents the embedding representation of a class. To generate these embeddings, MPNet [27] will be used. After obtaining the embeddings, we conduct a semantic search [28] using the cosine similarity to select the closest match among the COCO-Stuff classes. At the end of this stage, the Open-vocabulary classes with the predicted sizes will be ready to be used as input for the open-set object detector.
3.3 Open-set Object Detection and Segmentation
The input of this stage is the open-vocabulary set with the predicicted sizes that has been generated by the previous component, STS.
Object detection and segmentation are critical tasks in computer vision, particularly when dealing with unconstrained environments [21, 29, 30, 31]. Our method employs Grounding-DINO [20] as the object detector, querying it with curated textual labels for efficient semantic information extraction and object detection. Grounding-DINO generates bounding boxes, which are then used to prompt the Segment Anything Model (SAM) [13] to create high-quality instance masks for each detected object.
To further refine our approach, we employ Non-Maximum Suppression. During this stage, we address potential redundancies or errors introduced during object detection. Specifically, we focus on rectifying instances involving overlapping bounding boxes assigned identical labels or similar bounding box sizes but different labels, as shown in fig. 2. This approach distinct from per-class non-maximum suppression (NMS), which is commonly used in object detection [32]. In mathematical terms:
  1. 1.
[…]
  2. 2.
    In case of using the identified tags, when encounter bounding boxes that have equivalent areas but distinct class labels, the algorithm assigns the priority to the labels originating from captions rather than those derived from previously identified tags. As a result, only one bounding box will be used while others get removed through suppression.
    Mathematically, given two bounding boxes $b_{c}$ $b_{d}$ sharing same dimensions but disparate labels $y(b_{c})\neq y(b_{d})$, we prioritize retaining $b_{c}$ whenever its associated label stems from the caption. Otherwise, $b_{d}$ will be retained.
At the end of this stage, the detected objects and panoptic segmentation will be ready to be utilized by the next component, Unified Panoptic Mapping.
3.4 Unified Panoptic Mapping
The inputs of this component are the posed RGBD data, the open-vocabulary classes with their predicted sizes, the detected objects and panoptic segmentation.
In our study, we build upon the object-centric mapping framework introduced in [12]. So each submap contains all related data such as panoptic, instance, and class labels, as well as transformation and tracking data. Inspired by their semantically consistent representation of the environment, we adopt their notion of dividing the world into submaps for accurate temporal alignment and reduced computational complexity. However, we aim to improve the quality of objects contained within the maps. By integrating posed RGBD data with Unified Semantics, we ensure consistent object identification across different semantic labels through dynamic labeling, as illustrated in fig. 3. This approach simplifies classification and enhances the intuitive interaction with spatial information.
At the end of this stage, we will have a dynamically labeled multi-resolution multi-TSDF map.
Figure 3: Unified Semantics in Action: Three sequential frames featuring the same table with distinct semantic labels – ”table”, ”small dining table”, and ”wooden table”. Despite varied descriptions, the unified semantics maintains a consistent category ID across all instances, matching that of the COCO-Stuff parent class ”table”. This approach ensures semantic cohesion while allowing for more specific descriptors.
4 Experiments
To evaluate our proposed method, we compare UPPM performance with MaskDino [5] and the reference data segmentation provided by the evaluation dataset. MaskDINO is a closed-set model that recognizes a wide range of categories, namely LVIS1 [33] categories (1203 categories). Our goal is to demonstrate the effectiveness of UPPM with the dynamically generated customized categories in comparison with these strong baselines. We conduct our experiments on three indoor evaluation datasets, one of them is simulated and the others are real-world datasets.
The Flat dataset [12] as a simulated benchmark, showcasing its effectiveness in scene understanding and reconstruction. Despite its 31-category dataset segmentation, the Flat dataset provides comprehensive information, highlighting the challenges in achieving full scene understanding. We choose the Flat dataset to highlight the importance of detailed labels for improved scene reconstruction.
The RIO dataset has noisy conditions, which create challenges for our pipeline, particularly in terms of reducing error propagation between pipeline components (section 3).
The ScanNet v2 dataset [34] has more categories and higher image quality than the previous dataset, making it a more stringent evaluation benchmark.
4.1 Experimental Setup
We measure the performance using root-mean-square error (RMSE), mean-absolute error (MAE), Chamfer distance, and coverage. As shown in eq. 2, the Chamfer distance $d_{\text{ch}}(G,R)$ is computed by combining the sum of RMSE in both directions, $\text{RMSE}_{G\to R}$ and $\text{RMSE}_{R\to G}$, providing the errors from the ground truth point cloud $G$ to the evaluated point cloud $R$ and vice versa.
$d_{\text{ch}}(G,R)=\underbrace{\sum_{g\in G}\min_{r\in R}\left\|g-r\right\|^{2}_{2}}_{\text{RMSE}_{G\to R}}+\underbrace{\sum_{r\in R}\min_{g\in G}\left\|g-r\right\|^{2}_{2}}_{\text{RMSE}_{R\to G}}$ (2)
Root-mean-square error (RMSE) and mean-absolute error (MAE) are asymmetrical metrics, as the distance from ground truth points ($G$) to reconstructed map points ($R$) may differ from $R$ to $G$. These metrics compute distances by comparing each point in one set to its nearest neighbor in the other. A larger $\text{RMSE}_{R\to G}$ suggests potential inaccuracies in the reconstruction process, emphasizing the importance of accuracy for both sets.
As shown in eq. 3, the coverage is calculated as the percentage of observed ground truth points in the map:
$\text{Cov}=\frac{N_{\text{observed}}}{N_{\text{total}}}\times 100\%$ (3)
The experiments were conducted on a system with a 2.6 GHz CPU, an NVIDIA GeForce RTX 2060 6 GB GPU, and 16 GB of RAM, supplemented by a server featuring an NVIDIA GeForce RTX 2080 Ti 16 GB GPU and 32 GB of RAM.
4.2 Evaluations on the Flat dataset
In table 1, we present comparative quantitative results demonstrating the superior accuracy and competitive coverage of our UPPM model in comparison to both the dataset segmentation and the MaskDINO method [5].
Table 1: Quantitative comparison on the Flat dataset [12].
[…]
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
Dataset
Seg. [12]
1.27 2.24 0.66 0.80 3.05 71.30
MaskDINO
[5]
1.26 2.23 0.68 0.85 3.08 71.69
UPPM (Ours)
1.25 2.06 0.65 0.81 2.87 70.76
Figure 4: Qualitative Comparison of Map Reconstruction on Flat dataset [12].
It is worth noting that while MaskDINO achieves better coverage, a comprehensive comparison requires examining fig. 4. Upon inspection, it becomes apparent that MaskDINO has better reconstruction for structures like walls and floors. However, it fails to capture intricate details, like the paintings on the wall, which hold more semantic value. Conversely, UPPM successfully reconstructs these fine-grained elements.
4.3 Evaluations on ScanNet dataset
As demonstrated in table 2, UPPM exhibits better accuracy than MaskDINO [5] with a slight improvement in coverage. Although UPPM demonstrates strong performance, there is still room for improvement to surpass the accuracy scores achieved by the dataset segmentation [34].
Table 2: Comparative Evaluation Results on ScanNet v2 dataset [34].
[…]
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
Dataset
Seg. [12]
1.38 2.06 2.80 7.41 9.46 82.5
MaskDINO
[5]
1.81 2.51 4.40 14.26 16.76 81.21
UPPM (Ours)
1.78 2.43 3.30 10.21 12.64 81.42
4.4 Evaluations on the RIO dataset
The table 3 summarizes our quantitative findings, whereas fig. 5 shows our qualitative comparison. UPPM outperforms dataset segmentation [12] and MaskDINO [5] in accuracy while preserving good coverage.
Figure 5: Comparative Visualization. Top Row: Dataset segmentation with spatial distribution (Right) and matching reconstructed colored map (Left) for better scene understanding. Bottom Row: Our proposed UPPM demonstrates refined segmentation with pronounced clarity and detail in both spatial (Right) and colored map (Left) representations, highlighting the performance of UPPM in intricate scene understanding.
Table 3: Quantitative comparison on the RIO dataset [12].
[…]
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
Dataset
Seg. [12]
1.23 1.68 1.91 8.57 10.25 77.09
MaskDINO
[5]
1.28 1.74 1.97 7.42 9.16 76.63
UPPM (Ours)
1.19 1.63 1.55 4.02 5.65 74.14
5 Ablation Studies
The objective of these studies is to gain a deeper understanding of the contributions made by different components in our UPPM model.
5.1 Unified Semantics
In this study, we examine the effects of deactivating the unified semantic mechanism, which we call it PPM (Promptable Panoptic Mapping). In other words, we want to assess the impact of unified semantics on map reconstruction. As shown in tables 4, 5 and 6, PPM and UPPM have comparable performance across most metrics. Our UPPM approach, provides balanced performance across datasets, offering competitive accuracy, high coverage, and the added benefit of unified-semantic mechanism. As a result, UPPM offers a versatile alternative for applications where point-to-point precision and reconstruction completeness are equally important. As shown in fig. 6, PPM lacks dynamically labeled classes, leading to ambiguity when many semantic categories attach to the same object. In comparison, UPPM exhibits consistent behavior by reliably assigning semantic classes to objects while maintaining richness in dynamic labeling.
Figure 6: Qualitative Comparison of PPM vs UPPM in 3D-Semantic Segmentation. The predicted semantic classes are represented by $x,y,\&z$.
Table 4: Quantitative Ablation Experiments on the Flat dataset [12].
[…]
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
UPPM
1.25 2.06 0.65 0.81 2.87 70.76
UPPM with tags
1.25 2.01 0.64 0.78 2.80 70.85
PPM
1.21 1.99 0.65 0.80 2.80 70.38
PPM with tags
1.23 2.02 0.64 0.80 2.82 71.26
CPM
0.90 1.39 0.62 0.73 2.12 44.84
Table 5: Quantitative Ablation Experiments on the ScanNet v2 [34].
[…]
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
UPPM 1.78 2.43 3.30 10.21 12.64 81.42
UPPM with tags 1.74 2.41 3.66 11.21 13.63 82.64
PPM 1.79 2.48 3.06 9.17 11.64 81.07
PPM with tags 1.74 2.39 3.36 10.11 12.50 82.75
CPM 1.51 2.03 2.25 5.43 7.46 59.3
Table 6: Quantitative Ablation Experiments on the RIO data [12].
[…]
Method
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
MAE
[m]($\downarrow$)
RMSE
[m]($\downarrow$)
Chamfer
dist.
[m]($\downarrow$)
Cover
-age
[%]($\uparrow$)
UPPM 1.19 1.63 1.55 4.02 5.65 74.14
UPPM with tags 1.31 1.77 1.87 8.05 9.82 75.14
PPM 1.23 1.70 1.38 2.56 4.26 72.47
PPM with tags 1.23 1.70 1.73 7.30 9.00 73.49
CPM 1.11 1.46 2.76 14.55 16.01 43.42
5.2 Non-Maximum Suppression
To evaluate NMS (section 3.3) in our pipeline, we run the experiment on the Flat dataset. As shown in table 7, NMS improves coverage by $8.27\%$ gain, which proves that NMS widens the reconstructed area by eliminating overlapping or redundant predictions and enabling the pipeline to capture a wider range of discrete regions in the scene. This causes a small rise in Chamfer distance. The negligable decrease in point-to-point accuracy ($0.008[m]$ about $0.3\%$) can be attributed to the addition of less confident predictions in newly covered areas, and the occasional removal of valid overlapping predictions. In conclusion, NMS have significantly improved the overall performance.
Table 7: Quantitative results of NMS on the Flat dataset [12].
Method Chamfer dist. [m]($\downarrow$) Coverage [%]($\uparrow$)
PPM 2.798 70.38
PPM w/o NMS 2.790 62.11
Enhancement -0.008 [m] +8.27 %
5.3 Blurry Frames Filtering
Our work, as shown in fig. 2, uses the generated caption [18] to detect and skip the blurry frames, The impact of skipping blurry frames varies across the datasets:
Since the Flat dataset is a simulation dataset with high-quality images, blurry frames filtering has no effect on it.
For the ScanNet dataset, we notice that $0.84\%$ of the images are flagged as blurry, considerably less than those found in RIO. In the RIO dataset, we find that the blurry filter detects around $3\%$ of the data as being blurry. However, manual inspection and classification of the data revealed that more than $21\%$ of the dataset is affected by motion blur or other forms of degradation. This discrepancy suggests that the blurry filter identifies only extreme cases, leaving many less severe cases undetected. Our UPPM enhances the performance by reducing Chamfer distance by $16.675\%$ and increases coverage by $6.47\%$.
5.4 Identified Tags
As shown in fig. 2, we offer the option to use the identified tags in the pipeline. The effect of integrating the identified tags varied across datasets:
For the Flat dataset, as shown in the table 4, integrating the identified tags to both PPM and UPPM results in marginal improvement in coverage and all other metrics in the case of UPPM. This shows that semantic information adds to more comprehensive scene knowledge.
In both datasets, ScanNet and RIO, as shown in table 5 and table 6 respectively, adding the identified tags to both PPM and UPPM results in marginal improvement in coverage but, unlike the flat dataset, this addition increases Chamfer distance. Based on these results, we conclude that in instances when the dataset entails noise, depending exclusively on the UPPM and excluding the identified tags seems to offer better results. This maintains the advantages of dynamic labeling and unified semantics while preventing low-quality tags from corrupting the input to the open-set object detector.
5.5 Closed-Panoptic Mapping Performance
As part of the ablation studies, we replace all of the VLFE++, STS, and the Open-set Object Detector (check fig. 2) with a state-of-the-art closed-set object detector, namely Co-DINO [6]. We call this baseline Closed-Panoptic Mapping (CPM). The performance of CPM varied across datasets:
In Flat and ScanNet datasets, in terms of error measures, as shown in tables 4 and 5, the CPM baseline demonstrates better precision in point cloud reconstruction. However, it falls severely short in terms of coverage, indicating a trade-off between accuracy and completeness of reconstruction. From a qualitative perspective, as seen in fig. 7, CPM has difficulty identifying unforeseen objects during identification tasks.
In the RIO dataset, as shown in table 6, besides the low coverage, CPM exhibited the worst Chamfer distance. This behavior is attributed to substantial deviations from the ground truth during map reconstruction, leading to reduced accuracy and reliability.
Figure 7: Qualitative Comparison for CPM vs UPPM on Map Reconstruction
6 Conclusion
In this work, we propose UPPM, a novel approach that tackles the challenges of generating rich and accurate object labels for panoptic mapping by harnessing the power of dynamic labeling. Our system efficiently integrates diverse, possibly noisy labels from multiple sources into a consistent semantic structure, leading to effective postprocessing and precise segmentation. We demonstrated the usefulness of foundation models and their efficiency in being utilized in downstream mapping tasks without requiring model retraining.